OpenStudy (anonymous):

Which of the following would best represent a cosine function with an amplitude of 3, a period of pi over 2, and a midline at y = −4? f(x) = −4 cos 4x + 3 f(x) = 3 cos(x − pi over 2) − 4 f(x) = 4 cos(x − pi over 2) + 3 f(x) = 3 cos 4x − 4

OpenStudy (anonymous):

@LegendarySadist

OpenStudy (anonymous):

The amplitude is the number before the trig function (sine, cosine, etc.). So we want a 3 before the cos in this case.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

can u help

OpenStudy (anonymous):

It will usually be written out like this \[\large acos(bx-c)+d\\\large amplitude~=~a\\\\\large period~=~\frac{2pi}{b}\]

OpenStudy (anonymous):

The midline is the point halfway between the top and bottom of the wave. The midline of a cos function is usually 0. So to change it to 4 smaller, we have to make the vertical change (the d in my example) equal -4. So set d=-4

OpenStudy (anonymous):

A

OpenStudy (anonymous):

Wrong. Did you just guess?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

Nevermind i figured it out thanks anyways

OpenStudy (anonymous):

\[\large amplitude=a\\\large amplitude=3\\\large a=3\] So in \(\large acos(bx)+d\), \(\large a=3\)

OpenStudy (anonymous):

So did you get the answer then?